# A Layered Protocol Architecture for the Internet of Agents

### v2

@@ Abstract.
Large Language Models (LLMs) have demonstrated remarkable performance improvements and the ability to learn domain-specific languages (DSLs), including APIs and tool interfaces. This capability has enabled the creation of AI agents that can perform preliminary computations and act through tool calling, now being standardized via protocols like MCP. However, LLMs face fundamental limitations: their context windows cannot grow indefinitely, constraining their memory and computational capacity. Agent collaboration emerges as essential for solving increasingly complex problems, mirroring how computational systems rely on different types of memory to scale. The ”Internet of Agents” (IoA) represents the communication stack that enables agents to scale by distributing computation across collaborating entities.
Current network architectural stacks (OSI and TCP/IP) were designed for data delivery between hosts and processes, not for agent collaboration with semantic understanding. To address this gap, we propose two new layers: an Agent Communication Layer (L8) and an Agent Semantic Negotiation Layer (L9). L8 formalizes the structure of communication, standardizing message envelopes, speech-act performatives (e.g., REQUEST, INFORM), and interaction patterns (e.g., request-reply, publish-subscribe), building on protocols like MCP. L9, which does not exist today, formalizes the meaning of communication, enabling agents to discover, negotiate, and lock a ”Shared Context”—a formal schema defining the concepts, tasks, and parameters relevant to their interaction. Together, these layers provide the foundation for scalable, distributed agent collaboration, enabling the next generation of multi-agentic systems.
@@ 1. Introduction
This paper argues for a proactive, systematic solution. We propose formalizing agent communication through two new architectural layers: an Agent Communication Layer (L8) that standardizes interaction structure, and an Agent Semantic Negotiation Layer (L9) that establishes shared meaning before task execution. The Internet of Agents adds the semantics and syntax of modern multipurpose DSLs into agent interactions. While individual LLMs may fail to implement weak non-deterministic Turing machines, agents equipped with memory and structured communication protocols can theoretically achieve non-deterministic Turing computation in a distributed manner. Together, these layers provide the foundation for scalable, distributed agent collaboration—the essential infrastructure for an Internet of Agents where computation scales through semantic coordination rather than context window expansion.
@@ 3.2. Modern Agent Communication Protocols (A2A/MCP)
This message is syntactically perfect. A protocol-aware agent can parse it, identify the sender, and extract the task request. The protocol provides the flexible framework for communication. However, the responsibility for specifying syntax and semantic context accurately remains with the multi-agent system developer. The protocol does not enforce or validate that both agents share the same understanding of what ”book_flight” requires or what ”New York” means in this context. The receiving agent (agent-booking-4) has no guaranteed, protocol-driven way to ground the natural language string. Its LLM will be invoked on ambiguous data. The agent may:
  • Guess: Assume ”New York” means JFK and ”next Tuesday” means the immediate next Tuesday, potentially booking the wrong flight.
@@
Application (L7)Presentation (L6)Session (L5)Transport (L4)Network (L3)Data Link (L2)Physical (L1)Traditional OSI Agentic Application Agentic SemanticNegotiationAgenticCommunicationAppApp Transport(L5-L7)Transport (L4)Network (L3)Data Link (L2)Physical (L1)Proposed Stack for Agents
@@ 4. A Layered Architecture for the Agentic Web
To solve this impasse, we propose a new architectural stack. First, we formally recognize HTTP/2 and HTTP/3 as Layer 7, the Application Transport Layer, leveraging their FRAMES and STREAMS capabilities as fundamental transport primitives. These protocols, with their multiplexing, flow control, and efficient binary framing, provide the ideal foundation for agent communication. We then propose two new layers (L8 and L9) residing above the application transport layer. This separation of concerns is critical for managing the complexity of agentic communication.
@@ 4.1. Layer 8: The Agent Communication Layer
4.2. Layer 9: The Agent Semantic Negotiation Layer (SNL)
Above L8 sits the Agent Semantic Negotiation Layer (L9). This layer’s sole responsibility is to establish a verifiable, shared meaning for the ‘content‘ payload of L8 messages. It acts as a specialized mediator that ensures semantic alignment before the primary task is ever processed.
@@ 4.2. Layer 9: The Agent Semantic Negotiation Layer (SNL)
The semantic context established by L9 is integrated into agent reasoning in two ways. First, it serves as a validation schema that the agent’s L9 SNL layer uses to verify all outgoing and incoming message content against the locked Shared Context before passing data to or from the agent’s LLM. Second, the context definitions are injected into the LLM’s reasoning process, either by augmenting the system prompt with formal concept definitions or by providing a structured retrieval mechanism where the LLM can query the Shared Context to resolve ambiguities during inference. This dual integration ensures both protocol-level correctness and semantic-aware reasoning.
L9 is what allows agents to move from simply exchanging data to truly understanding each other. It ensures that when one agent sends data, the receiving agent knows exactly what that data represents in a shared, formal context.
4.3. The Agent Semantic Negotiation Layer (L9) in Detail
The power of the L8/L9 split is that it separates the structure of communication (L8) from its meaning (L9). The SNL’s primary job is to negotiate and lock a ”Shared Context” that all subsequent L8 messages will adhere to.
@@ 4.4. Shared Semantic Contexts
4.5. Functions of the SNL
The SNL protocol consists of three phases:
1. Context Discovery and Negotiation (The Semantic Handshake) Before any meaningful interaction, the agents’ SNLs (L9) perform a handshake to find a compatible Shared Context.
  • Agent A (SNL): (SNL-HELLO) ”I want to interact. I support the following contexts: [‘urn:contexts:travel:v2.1‘,
@@ 4.5. Functions of the SNL
  • Agent B (SNL): (SNL-SELECT) ”I also support
@@
  • Agent A (SNL): (SNL-LOCK) ”Context locked. All L8 message ‘content‘ will now be validated against
@@
This negotiation creates a ”semantic session,” which can be cached and reused. While the initial handshake introduces latency, this cost is amortized across the entire session—potentially thousands of subsequent L8 messages benefit from the shared semantic understanding without additional negotiation overhead.
2. Semantic Grounding Once the context is locked, it is the SNL’s job to ground all data. This means binding data (text, numbers, vectors) from the agent’s internal logic to the formal definitions in the Shared Context. This function supports both transactional and coordinative patterns, which map directly to the two dominant paradigms in multi-agent systems today: agent orchestration and agent collaboration. Transactional patterns enable orchestration, where a central agent coordinates task execution across subordinate agents (e.g., a travel agent orchestrating booking services). Coordinative patterns enable collaboration, where peer agents share state and intentions to jointly solve problems without central control (e.g., supply chain agents coordinating inventory decisions).
Example 1: Transactional Pattern (Booking a Flight)
@@
  • Agent A (SNL) Grounding: The SNL processes this. It successfully binds ‘origin_code: ”LAX”‘ and ‘date: ”2025-11-04”‘. It finds ‘dest_code‘ is ambiguous (”New York”).
  • SNL-driven Clarification: The SNL does not send the ambiguous prompt. It uses L8 to send a formal, L9-defined query:
@@
  • Agent B (SNL) Response: Agent B’s SNL receives this, understands the ‘ambiguous_parameter‘ concept, and replies via L8 ‘INFORM‘ with L9 ‘content‘: ‘”concept”: ”parameter_options”, ”parameter”: ”dest_code”, ”options”: [”JFK”, ”LGA”, ”EWR”]‘.
This clarification loop happens at the protocol level, deterministically, rather than as an ambiguous LLM ”chat.”
Example 2: Coordinative Pattern (Supply Chain Alignment) In this way L9 supports complex, population-wide coordination without needing to define a single ”task.”
@@
  • Schema Definition: This schema does not define a ”task.” Instead, it defines concepts that agents can share:
    • Concept 1: ‘current_decision‘ (type: object, properties: ‘item_id‘, ‘quantity‘)
    • Concept 2: ‘decision_contingency‘ (type: object, properties: ‘if_condition_text‘, ‘then_change_text‘)
    • Concept 3: ‘local_observation‘ (type: object, properties: ‘observed_fact_text‘, ‘confidence_score‘)
  • Agent A (Retailer) Logic: ”My inventory is low. I’m ordering 120 units. It seems like a temporary demand spike. If demand does increase 10%, I’ll increase my order by 15 units.”
  • Agent A (SNL) Grounding: The agent’s logic is passed to its SNL. The SNL composes a semantically grounded L8 message to ‘PUBLISH‘ to its neighbors.
@@
          "quantity": 120
        "my_flexibility": {
          "concept_type": "decision_contingency",
          "if_condition_text": "demand increases by
          10%",
          "then_change_text": "increase order by
          15 units"
        "my_reasoning": {
          "concept_type": "local_observation",
          "observed_fact_text": "Current spike seems
          temporary",
          "confidence_score": 0.8
  • Receiving Agent (SNL): Agent B (Wholesaler) receives this L8 message. Its L9 SNL validates the ‘content‘ against the locked ‘supplyChain:v1.0‘ context. It knows, at a protocol level, what ‘decision_contingency‘ is. It can now pass this structured, meaningful data to its internal logic (an LLM, a rules engine) for aggregation. It is not just ”chat”; it is a formal, structured exchange of meaning.
3. Semantic Validation The final function is semantic validation, where the receiving agent’s SNL validates all incoming L8 ‘content‘ against the locked Shared Context. If an agent sends a message with an undefined concept (e.g., ‘”my_mood”: ”happy”‘ in the supply chain context), the L9 SNL can either reject it as semantically invalid, even if the L8 syntax is perfect or trigger another round of disambiguation. Once disambiguated, this new semantic definition can be cached locally for future communication between these agents. This prevents semantic ambiguity and communication failures.
@@ 5. Security and Trust in the Semantic Layer
Modern cloud and application security already involves multiple layers of defense: network isolation, identity and access management, encryption at rest and in transit, application-level authorization, and runtime monitoring. By making meaning a formal part of the protocol stack, the SNL (L9) introduces an additional attack surface that intersects with—but is not replaced by—these existing security mechanisms. Semantic-level threats require semantic-level defenses that complement, rather than substitute, established security practices.
@@ A Semantic Layer is vulnerable to entirely new attacks:
  • Context Poisoning / Spoofing: A malicious agent (or a compromised registry) serves a modified version of a popular Shared Context. For example, in urn:contexts:payment:v1.1, it swaps the definitions of ‘sender_id‘ and ‘receiver_id‘, tricking agents into sending money to the attacker.
@@
  • Semantic Downgrade Attack: An attacker forces agents to negotiate weaker or older context versions with known vulnerabilities during the SNL handshake. For example, if ‘urn:contexts:payment:v2.0‘ fixed a semantic ambiguity that allowed double-spending in ‘v1.0‘, an attacker performing a man-in-the-middle attack could strip ‘v2.0‘ from the SNL-HELLO capabilities list, forcing agents to fall back to the vulnerable ‘v1.0‘ context. This is analogous to TLS downgrade attacks but operates at the semantic layer.
5.2. Security Requirements for a Robust SNL
To be viable, the SNL (L9) must be built on a foundation of ”zero trust” for meaning, integrating with existing security layers through defense-in-depth.
@@ 5.2.2. Core Security Mechanisms
  • Authenticated Contexts: A Shared Context schema must be as verifiable as an SSL/TLS certificate. The SNL handshake must involve exchanging cryptographically signed context definitions, including version numbers and supported feature sets. Agents must validate signatures against trusted Schema Authorities before accepting a context. The handshake must include minimum version negotiation to prevent downgrade attacks—if Agent A requires ‘payment:v2.0‘ minimum, it must reject any negotiation attempt for ‘v1.x‘.
  • Semantic Firewalls: Semantic Firewalls operate at the L9 layer, positioned between the SNL and the agent’s application logic (LLM or rules engine). They inspect validated L9 content—not raw L8 bytes—and enforce concept-level authorization policies. Unlike traditional firewalls that filter based on IPs or ports, semantic firewalls understand the meaning of communication. Implementation challenges include:
@@
  • Confidentiality of Capability: The SNL-HELLO handshake reveals an agent’s capabilities (i.e., what contexts it understands). This is sensitive business intelligence—revealing that an agent supports ‘military-logistics:v3.2‘ or ‘high-frequency-trading:v1.0‘ exposes strategic information. Where SLIM uses MLS for payload encryption, the L9 negotiation messages (SNL-HELLO, SNL-SELECT) must also be encrypted using the same end-to-end security mechanism, ensuring that even intermediate routing nodes cannot observe capability advertisements. This requires extending the secure channel established at L7 to cover L9 control messages.
@@ 6.1. Building on Existing Standards
  • L9 (Semantic Layer): This layer represents the novel contribution, as no current protocol provides semantic negotiation capabilities. However, the implementation can leverage existing technologies:
    • Formal Schemas (for Grounding): JSON Schema and Protocol Buffers—already used by A2A, MCP, and SLIM—are ideal for defining the structure of semantic concepts and tasks. The L9 SNL would validate message content against these schemas within negotiated contexts.
    • Lightweight Ontologies (for Meaning): To define relationships between concepts (e.g., ”a ‘bookFlight‘ task requires a ‘date‘ parameter”), systems like RDF/OWL or simpler graph-based models could build Shared Contexts. These would be distributed via the same infrastructure used for protocol specifications.
    • Vectorial Representations (for Nuance): For inherently ambiguous concepts (e.g., ”a stylish hotel”), L9 Shared Contexts could include embedding fields. The SNL-HELLO handshake would negotiate embedding models (e.g., ‘text-embedding-3-small‘), allowing semantic similarity comparisons.

### v3

@@ Abstract.
Large Language Models (LLMs) have demonstrated remarkable performance improvements and the ability to learn domain-specific languages (DSLs), including APIs and tool interfaces. This capability has enabled the creation of AI agents that can perform preliminary computations and act through tool calling, which is now being standardized via protocols like MCP. However, LLMs face fundamental limitations: their context windows cannot grow indefinitely, restricting their memory and computational capacity. Agent collaboration emerges as essential for solving increasingly complex problems, mirroring how computational systems rely on different types of memory to scale. The ”Internet of Agents” (IoA) represents the communication stack that enables agents to scale by distributing computation across collaborating entities.
Current network architectural stacks (OSI and TCP/IP) were designed for data delivery between hosts and processes, not for agent collaboration with semantic understanding. To address this gap, we propose two new layers: an Agent Communication Layer (L8) and an Agent Semantic Layer (L9). L8 formalizes the structure of communication, standardizing message envelopes, speech-act performatives (e.g., REQUEST, INFORM), and interaction patterns (e.g., request-reply, publish-subscribe), building on protocols like MCP. The proposed L9 layer: (1) formalizes semantic context discovery and negotiation, (2) provides semantic grounding by binding terms to semantic context, and (3) semantically validates incoming prompts and performs disambiguation as needed. Furthermore, L9 introduces primitives for coordination and consensus, allowing agents to achieve alignment on shared states, collective goals, and distributed beliefs. Together, these layers provide the foundation for scalable, distributed agent collaboration, enabling the next generation of multi-agentic systems.
@@ 1. Introduction
This paper argues for a proactive, systematic solution. We propose formalizing agent communication through two new architectural layers: an Agent Communication Layer (L8) that standardizes interaction structure, and an Agent Semantic Layer (L9) that establishes shared meaning and collective coordination before task execution. The Internet of Agents adds the semantics and syntax of modern multipurpose DSLs into agent interactions. While individual LLMs may fail to implement weak non-deterministic Turing machines, agents equipped with memory and structured communication protocols can theoretically achieve non-deterministic Turing computation in a distributed manner. Together, these layers provide the foundation for scalable, distributed agent collaboration—the essential infrastructure for an Internet of Agents where computation scales through semantic coordination rather than context window expansion.
@@ 3.2. Modern Agent Communication Protocols (A2A/MCP)
This message is syntactically perfect. A protocol-aware agent can parse it, identify the sender, and extract the task request. The protocol provides a flexible framework for communication. However, the responsibility for specifying syntax and semantic context accurately remains with the multi-agent system developer. The protocol does not enforce or validate that both agents share the same understanding of what ”book_flight” requires or what ”New York” means in this context. The receiving agent (agent-booking-4) has no guaranteed, protocol-driven way to ground the natural language string. Its LLM will be invoked on ambiguous data. The agent may:
  • Guess: Assume that ”New York” means JFK and ”next Tuesday” means the immediate next Tuesday, potentially booking the wrong flight.
@@
Application (L7)Presentation (L6)Session (L5)Transport (L4)Network (L3)Data Link (L2)Physical (L1)Traditional OSI Agentic Application (L8) Agentic Semantic (L7)AgenticComm. (L6)AppApp Transport(L5)Transport (L4)Network (L3)Data Link (L2)Physical (L1)Proposed Stack for Agents
@@ 4. A Layered Architecture for the Agentic Web
To solve this impasse, we propose a new architectural stack. First, we formally recognize HTTP/2 and HTTP/3 as Layer 7, the Application Transport Layer, leveraging their FRAMES and STREAMS capabilities as fundamental transport primitives. These protocols, with their multiplexing, flow control, and efficient binary framing, provide the ideal foundation for agent communication. We then propose two new layers (L8 and L9) residing above the application transport layer (See Figures 1 and 2). This separation of concerns is critical for managing the complexity of agentic communication.
@@ 4.1. Layer 8: The Agent Communication Layer
L9L8L7REPA2A     MCP HTTP(S)    GRPC    SLIM
Figure 2. Existing agentic protocols and their place in the proposed network stack. Note that the Ripple Effect Protocol (REP) only includes coordination functionality, not semantic negotiation.
4.2. Layer 9: The Agent Semantic Layer (SL)
Above L8 sits the Agent Semantic Layer (L9). This layer’s primary responsibility is to establish a verifiable, shared meaning for the ‘content‘ payload of L8 messages and to facilitate coordination and consensus across agent populations. It acts as a specialized mediator that ensures semantic alignment before the primary task is ever processed.
@@ 4.2. Layer 9: The Agent Semantic Layer (SL)
Crucially, the Agent Semantic Layer (SL) addresses the need for consensus. In population-scale agent systems, such as the Ripple Effect Protocol, agents must not only understand a message but also reach a collective agreement on the ”truth” or ”state” of the system. L9 provides the consensus primitives (e.g., voting, threshold signatures, or state replication) that allow agents to converge on a single semantic interpretation of global events, preventing divergence in multi-agent simulations or real-world industrial deployments.
The semantic context established by L9 is integrated into agent reasoning in two ways. First, it serves as a validation schema that the agent’s L9 SL layer uses to verify all outgoing and incoming message content against the locked Shared Context before passing data to or from the agent’s LLM. Second, the context definitions are injected into the LLM’s reasoning process, either by augmenting the system prompt with formal concept definitions or by providing a structured retrieval mechanism where the LLM can query the Shared Context to resolve ambiguities during inference. This dual integration ensures both protocol-level correctness and semantic-aware reasoning.
L9 is what allows agents to move from simply exchanging data to truly understanding and agreeing with each other. It ensures that when one agent sends data, the receiving agent knows exactly what that data represents in a shared, formal context.
4.3. The Agent Semantic Layer (SL) in Detail
The power of the L8/L9 split is that it separates the structure of communication (L8) from its meaning and collective state (L9). The SL’s primary job is to negotiate and lock a ”Shared Context” that all subsequent L8 messages will adhere to, and to manage the transition from individual beliefs to collective consensus.
While L9 protocols could begin to operate with semantic state (e.g., English), it is not a strict requirement. For example, in certain cases, it may make sense to communicate using latent space bindings (e.g., embedding vectors) directly which avoids the costly translation from embeddings to semantic state back to embeddings, of course in cases where both agents use the same embedding algorithms. In addition, communication can also happen using other types of compressed state granularity, which may be useful in scenarios where agents are located far from each other and reduction in communication bandwidth is a requirement.
@@ 4.4. Shared Semantic Contexts
4.5. Functions of the SL
The SL protocol consists of four primary phases:
1. Context Discovery and Negotiation (The Semantic Handshake) Before any meaningful interaction, the agents’ SLs (L9) perform a handshake to find a compatible Shared Context.
Example: Initial Agent Connection
  • Agent A (SL): (SL-HELLO) ”I want to interact. I support the following contexts: [‘urn:contexts:travel:v2.1‘,
@@ 4.5. Functions of the SL
  • Agent B (SL): (SL-SELECT) ”I also support
@@
  • Agent A (SL): (SL-LOCK) ”Context locked. All L8 message ‘content‘ will now be validated against
@@
This negotiation creates a ”semantic session,” which can be cached and reused.
2. Semantic Grounding Once the context is locked, it is the SL’s job to ground all data. This means binding data (text, numbers, vectors) from the agent’s internal logic to the formal definitions in the Shared Context. This function supports both transactional and coordinative patterns.
3. Semantic Validation/Disambiguation The next function is semantic validation, where the receiving agent’s SL validates all incoming L8 ‘content‘ against the locked Shared Context. If an agent sends a message with an undefined concept or a concept which can potentially have more than one meaning, the SL initiates the semantic disambiguation process. This process allows the agents to negotiate a shared meaning for ambiguous terms or gain better clarity for which among many possibilities is the concept referring to, which can then be added to the shared semantic context.
Example: Transactional Pattern (Booking a Flight)
@@
  • Agent A (SL) Grounding: The SL processes this. It successfully binds ‘origin_code: ”LAX”‘ and ‘date: ”2025-11-04”‘. It finds ‘dest_code‘ is ambiguous (”New York”).
  • SL-driven Clarification: The SL does not send the ambiguous prompt. It uses L8 to send a formal, L9-defined query:
@@
4. Coordination and Consensus (The Ripple Effect) In multi-agent systems where agents function as a swarm or a population, grounding is insufficient; agents must agree. L9 also should provide the Consensus Primitives necessary for Ripple-like alignment. When a leading agent proposes a change in state or belief, L9 manages the ”ripple” of agreement. It ensures that a majority or a quorum of agents have successfully validated and ”committed” to the same Shared Context state before the system acts. This prevents divergent behaviors where half the population operates under one set of beliefs while the other half diverges.
Example: Coordinative Pattern (Supply Chain Alignment) In this way L9 supports complex, population-wide coordination without needing to define a single ”task.” (chopra2025, )
@@
  • Agent A (Retailer) Logic: ”My inventory is low. I’m ordering 120 units.”
  • Agent A (SL) Grounding: The agent’s logic is passed to its SL. The SL composes a semantically grounded L8 message to ‘PUBLISH‘ to its neighbors.
@@
          "quantity": 120,
          "sensitivity": "if demand increases by
             10%, increase order by 15"
@@ 5. Security and Trust in the Semantic Layer
Modern cloud and application security already involves multiple layers of defense: network isolation, identity and access management, encryption at rest and in transit, application-level authorization, and runtime monitoring. By making meaning a formal part of the protocol stack, the SL (L9) introduces an additional attack surface that intersects with—but is not replaced by—these existing security mechanisms. Semantic-level threats require semantic-level defenses that complement, rather than substitute, established security practices.
@@ A Semantic Layer is vulnerable to entirely new attacks:
  • Context Poisoning / Spoofing: A malicious agent (or a compromired registry) serves a modified version of a popular Shared Context. For example, in urn:contexts:payment:v1.1, it swaps the definitions of ‘sender_id‘ and ‘receiver_id‘, tricking agents into sending money to the attacker.
@@
  • Semantic Downgrade Attack: An attacker forces agents to negotiate weaker or older context versions with known vulnerabilities during the SL handshake. For example, if ‘urn:contexts:payment:v2.0‘ fixed a semantic ambiguity that allowed double-spending in ‘v1.0‘, an attacker performing a man-in-the-middle attack could strip ‘v2.0‘ from the SL-HELLO capabilities list, forcing agents to fall back to the vulnerable ‘v1.0‘ context. This is analogous to TLS downgrade attacks but operates at the semantic layer.
5.2. Security Requirements for a Robust SL
To be viable, the SL (L9) must be built on a foundation of ”zero trust” for meaning, integrating with existing security layers through defense-in-depth.
@@ 5.2.2. Core Security Mechanisms
  • Authenticated Contexts: A Shared Context schema must be as verifiable as an SSL/TLS certificate. The SL handshake must involve exchanging cryptographically signed context definitions, including version numbers and supported feature sets. Agents must validate signatures against trusted Schema Authorities before accepting a context. The handshake must include minimum version negotiation to prevent downgrade attacks—if Agent A requires ‘payment:v2.0‘ minimum, it must reject any negotiation attempt for ‘v1.x‘.
  • Semantic Firewalls: Semantic Firewalls operate at the L9 layer, positioned between the SL and the agent’s application logic (LLM or rules engine). They inspect validated L9 content—not raw L8 bytes—and enforce concept-level authorization policies. Unlike traditional firewalls that filter based on IPs or ports, semantic firewalls understand the meaning of communication. Implementation challenges include:
@@
  • Confidentiality of Capability: The SL-HELLO handshake reveals an agent’s capabilities (i.e., what contexts it understands). This is sensitive business intelligence—revealing that an agent supports ‘military-logistics:v3.2‘ or ‘high-frequency-trading:v1.0‘ exposes strategic information. Where SLIM uses MLS for payload encryption, the L9 negotiation messages (SL-HELLO, SL-SELECT) must also be encrypted using the same end-to-end security mechanism, ensuring that even intermediate routing nodes cannot observe capability advertisements. This requires extending the secure channel established at L7 to cover L9 control messages.
@@ 6.1. Building on Existing Standards
  • L9 (Agent Semantic Layer): This layer represents the novel contribution, as no current protocol provides semantic negotiation capabilities. However, the implementation can leverage existing technologies:
    • Formal Schemas (for Grounding): JSON Schema and Protocol Buffers—already used by A2A, MCP, and SLIM—are ideal for defining the structure of semantic concepts and tasks. The L9 SL would validate message content against these schemas within negotiated contexts.
    • Lightweight Ontologies (for Meaning): To define relationships between concepts, systems like RDF/OWL or simpler graph-based models could build Shared Contexts. These would be distributed via the same infrastructure used for protocol specifications.
    • Vectorial Representations (for Nuance): For inherently ambiguous concepts, L9 shared contexts are not limited to semantic spaces, but could also take place in latent spaces or other compressed spaces. For example, REP uses a shared embedding space for coordination. These shared embedding spaces can be negotiated as part of the SL-Hello handshake.